OpenStudy (anonymous):

Which represents the type of sequence: 12, 22, 30, 36, 41, …?

OpenStudy (anonymous):

both arithmetic and geometric

OpenStudy (anonymous):

shall i tell the answer ;) ?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

T2 - T1 = 10 T3 - T2 = 8 T4 - T3 = 6 Which tells its not arithmetic. T2 / T1 = 11/6 T3 / T2 = 15/11 T4 / T3 = 6/5 Which tells its not geometric. Hence, Neither arithmetic, nor geometric :P Happy happy? :D